jhonyy9 (jhonyy9):

opinion please ... One possible Proof of Collatz’s Conjecture Andrew John Gábor Gizella út 6/b 1143 Budapest - Hungary gbrndrs1968@yahoo.com ,,...Take any natural number n. If n is even, divide it by 2 to get n / 2. If n is odd, multiply it by 3 and add 1 to obtain 3n + 1....

jhonyy9 (jhonyy9):

One possible Proof of Collatz’s Conjecture Andrew John Gábor Gizella út 6/b 1143 Budapest - Hungary gbrndrs1968@yahoo.com ,,...Take any natural number n. If n is even, divide it by 2 to get n / 2. If n is odd, multiply it by 3 and add 1 to obtain 3n + 1. Repeat the process (which has been called "Half Or Triple Plus One", or HOTPO) indefinitely. The conjecture is that no matter what number you start with, you will always eventually reach 1. The property has also been called oneness. Paul Erdős said, allegedly, about the Collatz conjecture: "Mathematics is not yet ripe for such problems" and also offered $500 for its solution....” n1=1 --- 3x1+1 -- 4/2 -- 2/2 = 1 n2=2 --- 2/2 = 1 n3=3 --- 3x3+1 = 10/2 -- 5x3+1 -- 16/2 -- 8/2 -- 4/2 -- 2/2 = 1 n4=4 --- 4/2 -- 2/2 = 1 n5=5 --- 5x3+1 -- 16/2 -- 8/2 -- 4/2 -- 2/2 = 1 n6=6 --- 6/2 -- 3x3+1 -- 10/2 -- 5x3+1 -- 16/2 -- 8/2 -- 4/2 -- 2/2 = 1 n7=7 --- 7x3+1 -- 22/2 -- 11x3+1 -- 34/2 -- 17x3+1 -- 52/2 -- 26/2 -- 13x3+1 -- 40/2 -- 20/2 -- 10/2 -- 5x3+1 -- 16/2 -- 8/2 -- 4/2 -- 2/2 = 1 n8=8 --- 8/2 -- 4/2 -- 2/2 = 1 n9=9 --- 9x3+1 -- 28/2 -- 14/2 -- 7x3+1 -- 22/2 -- 11x3+1 -- 34/2 -- 17x3+1 -- 52/2 -- 26/2 -- 13x3+1 -- 40/2 -- 20/2 -- 10/2 -- 5x3+1 -- 16/2 -- 8/2 -- 4/2 -- -- 2/2 = 1 n10=10 --- 10/2 -- 5x3+1 -- 16/2 -- 8/2 -- 4/2 -- 2/2 = 1 …......................................................................................................................... 2. for n=(n2)n (n2)n=2n --- ... -- 5n2/2 -- (2n2 + 1)x3 + 1 -- 6n2 + 4 -- 8n2/2 -- 4n2/2 -- 2n2/2 -- 2/2 = 1 (n2)n+1=2n+1 --- ... -- 5n2/2 -- (2n2 + 1)x3 + 1 -- 6n2 + 4 -- 8n2/2 -- 4n2/2 -- 2n2/2 -- 2/2 = 1 for n = k suppose that is true (n2)n=(k2)n=2k --- ... -- 5k2/2 -- (2k2 + 1)x3 + 1 -- 6k2 + 4 -- 8k2/2 -- 4k2/2 -- 2k2/2 -- 2/2 = 1 (n2)n+1=(k2)n+1=2k+1 --- ... -- 5k2/2 -- (2k2 + 1)x3 + 1 -- 6k2 + 2k2 -- 8k2/2 -- 4k2/2 -- 2k2/2 -- 2/2 = 1 for k = k + 1 k2n=(k2n +1) --- … -- (5(k2+1))x3 +1 -- 15k2+15 +1 -- 15k2+16 -- 15k2+8k2 -- 23k2/2 -- (11k2+1)x3 +1 -- 33k2+2k2-- 35k2/2 -- (17k2+1)x3 +1 -- 51k2+ 2k2 -- 53k2/2 -- (26k2+1)x3 +1-- 78k2+2k2-- 80k2/2 -- 40k2/2 -- 20k2/2 -- 10k2/2 -- 5k2/2 -- (2k2+1)x3 +1 -- 6k2+2k2 -- 8k2/2 -- 4k2/2 -- 2k2/2 -- 2/2 = 1 k2n+1=(k2n+1 +1) --- … -- (5(k2 +1) +1) -- 5k2 +6 -- 5k2 +3k2 -- 8k2/2 -- 4k2/2 -- 2k2/2 -- 2/2 = 1 q.e.d.

OpenStudy (anonymous):

Together we can rule the galaxy!

OpenStudy (experimentx):

interesting

OpenStudy (experimentx):

what is for n=(n2)n ??

jhonyy9 (jhonyy9):

so n2 like a note wann being equal 2

jhonyy9 (jhonyy9):

so this mean 2n

OpenStudy (experimentx):

sorry, can't undertand ... could you use latex??

jhonyy9 (jhonyy9):

n2 and k2 mean 2 ok ?

OpenStudy (experimentx):

looks some people here might be interested in your work http://mathoverflow.net/

jhonyy9 (jhonyy9):

thank you

Directrix (directrix):

@jhonyy9 Do you have an Erdős number? If not, you should.

OpenStudy (anonymous):

too far above my head.

OpenStudy (anonymous):

like reading Laos' language

sam (.sam.):

You can post your findings here too, http://math.stackexchange.com/ , most of the advanced or professionals use it

jhonyy9 (jhonyy9):

yes i have posted it there and i have got for 1 reputation and i bronz badge with title of teacher thank you for your idea and opinion

OpenStudy (anonymous):

see in abc ab=ac ap=bc a=20 to find pbc

OpenStudy (anonymous):



OpenStudy (anonymous):

to find x ie pbc

OpenStudy (anonymous):

angle b is 80

Parth (parthkohli):

Is that a proof by cases?

jhonyy9 (jhonyy9):

do you see that just for cases ? by mathinduction we get for these cases from what we can deducting that is true for all numbers or not ?